OpenStudy (sonusuja):

(1+cotA+tanA)(sinA-cosA)=secA/sec^2A-cosecA/sce^2A

OpenStudy (learner):

\(\color{black}{\displaystyle \left(1+\cot A + \tan A\right)\left(\sin A - \cos A\right)=\frac{\sec A}{\sec^2A}-\frac{\csc A}{\csc^2A}}\) like this?

OpenStudy (learner):

or like that? \(\color{black}{\displaystyle \left(1+\cot A + \tan A\right)\left(\sin A - \cos A\right)=\frac{\sec A}{\sec^2A}-\frac{\csc A}{\sec^2A}}\)

OpenStudy (learner):

(I'm just trying to make sure my interpretation of the equation is correct, before we go ahead and work it out ... Are you there?)

jhonyy9 (jhonyy9):

i think the second have wrote it right ?